OpenStudy (anonymous):

Find the geometric mean of the pair of numbers 6 and 10

OpenStudy (anonymous):

is it 60?

hartnn (hartnn):

the geometric mean of a and b is \(\sqrt{ab}\)

OpenStudy (anonymous):

soooo... um that = 66?

hartnn (hartnn):

lol ,how ??

OpenStudy (anonymous):

so is it 60?

hartnn (hartnn):

ok, to find geometric mean, u first find product and then take its square root. u found the product correctly as 60 now take its square root.

OpenStudy (anonymous):

7.7?

hartnn (hartnn):

yes, u can keep it as \(\sqrt{60}= 2\sqrt{15}\) ok?

OpenStudy (anonymous):

oh ok